# PhysGM: Large Physical Gaussian Model for Feed-Forward 4D Synthesis

### v1

@@ Abstract
While physics-grounded 3D motion synthesis has seen significant progress, current methods face critical limitations. They typically rely on pre-reconstructed 3D Gaussian Splatting (3DGS) representations, while physics integration depends on either inflexible, manually defined physical attributes or unstable, optimization-heavy guidance from video models. To overcome these challenges, we introduce PhysGM, a feed-forward framework that jointly predicts a 3D Gaussian representation and its physical properties from a single image, enabling immediate, physical simulation and high-fidelity 4D rendering. We first establish a base model by jointly optimizing for Gaussian reconstruction and probabilistic physics prediction. The model is then refined with physically plausible reference videos to enhance both rendering fidelity and physics prediction accuracy. We adopt the Direct Preference Optimization (DPO) to align its simulations with reference videos, circumventing Score Distillation Sampling (SDS) optimization which needs back-propagating gradients through the complex differentiable simulation and rasterization. To facilitate the training, we introduce a new dataset PhysAssets of over 24,000 3D assets, annotated with physical properties and corresponding guiding videos. Experimental results demonstrate that our method effectively generates high-fidelity 4D simulations from a single image in one minute. This represents a significant speedup over prior works while delivering realistic rendering results. Our project page is at:https://hihixiaolv.github.io/PhysGM.github.io/
@@ Introduction
Recent advances in 3D representation, particularly 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023), have revolutionized novel-view synthesis for static scenes. The next frontier is to imbue these representations with dynamic, physically plausible behavior, unlocking applications in virtual reality, robotics (Lu et al. 2024), and autonomous systems (Zhou et al. 2024). However, creating high-fidelity, physics-based 4D content remains a significant challenge, often demanding computationally expensive, per-scene optimization pipelines that are ill-suited for real-time or large-scale deployment.
@@
Current paradigms for physics-based 4D synthesis are hampered by a fundamental bottleneck: a reliance on slow, iterative optimization. The typical workflow involves first reconstructing a 3DGS model from dense multi-view images, then manually specifying its physical properties (e.g., stiffness, mass) via configuration files (Xie et al. 2024), and finally running a physics simulation. This process is not only computationally prohibitive but also lacks scalability and generalization. While recent works have explored learning physical properties by distilling knowledge from video models (Lin et al. 2025; Zhang et al. 2024c), these methods still require extensive per-scene training and fail to address the core issue of efficiency.
@@
To overcome these limitations, we ask a fundamental question: Can we bypass per-scene optimization entirely and instead learn a generative model that produces a complete, physically-grounded 4D simulation in a single, feed-forward pass? This requires reframing the problem from one of slow, iterative reconstruction to one of amortized, feed-forward inference. To this end, we introduce PhysGM, a physics-grounded generative model that enables optimization-free synthesis of dynamic 4D scenes from sparse image inputs. Our key insight is a novel two-stage training paradigm designed to learn a generalizable physical prior and for perceptual realism. In the first stage, we pre-train our model on a substantial dataset to jointly predict a 3D Gaussian representation and its corresponding physical properties. This establishes a robust generative prior. In the second stage, we employ Direct Preference Optimization (DPO) (Rafailov et al. 2023) to fine-tune the model. By ranking generated simulations against a ground-truth video, we create preference pairs that guide the model towards producing more physically plausible and temporally coherent dynamic sequences, all without requiring a differentiable physics engine.
@@
To facilitate this research, we construct and release the PhysAssets Dataset, a substantial benchmark of over 24,000 3D assets, each annotated with its physical material properties and a corresponding physically-plausible simulation video, providing a critical resource for training and evaluating generative 4D models.
@@
  • We introduce PhysGM, the first feed-forward framework capable of generating a complete, physically-grounded 4D Gaussian simulation from sparse image inputs in under a minute.
@@
  • We propose a novel two-stage training paradigm that combines large-scale supervised pre-training with DPO-based refinement, enabling the model to learn a robust physical prior and then align it with perceptual quality.
@@
  • We present the PhysAssets Dataset, a new substantial benchmark with 24,000+ assets annotated with physical properties and simulation videos, to spur future research in this domain.
Related Work
4D Content Generation
The generation of 4D content—dynamic 3D scenes with temporal evolution—has recently seen rapid progress, largely branching into two main paradigms: distillation from 2D generative models and direct supervision on 3D data.
The first paradigm leverages powerful pre-trained 2D or video diffusion models as priors, using Score Distillation Sampling (SDS)(Poole et al. 2022) to optimize a dynamic 3D representation(Yin et al. 2023; Ren et al. 2023; Bahmani et al. 2024; Ling et al. 2024). Some approaches explicitly model temporal deformations, for instance, by using warped Gaussian surfaces (Wang et al. 2023a). To improve temporal consistency, other works employ multi-view video models for more robust supervision (Zhang et al. 2024a; Jiang et al. 2024). However, the iterative nature of SDS is computationally expensive and can be unstable. To mitigate this, some methods pre-synthesize full video sequences to use as direct supervision (Ren et al. 2023; Pan et al. 2024), while others propose generalizable models to bypass per-scene optimization entirely (Ren et al. 2024). The second paradigm trains generative models directly on paired text-3D data (Jun and Nichol 2023; Nichol et al. 2022). Pioneering works like Point-E (Nichol et al. 2022) generated point clouds, while more recent models like Shap-E (Jun and Nichol 2023) and LGM (Tang et al. 2024) produce implicit functions (SDFs)(Park et al. 2019) and explicit 3DGS representations(Kerbl et al. 2023), respectively. The primary drawback of this approach is its reliance on large-scale 3D datasets, which remain scarce.
@@
Despite their visual progress, these methods are fundamentally agnostic to the laws of physics. Their focus on visual plausibility often yields animations that are temporally coherent but physically impossible, as they lack an explicit model of real-world dynamics.
Feed-forward 3D Gaussian Models
@@
Parallel to 4D generation, another line of research has focused on accelerating the creation of static 3D scenes. Traditional 3DGS reconstruction (Kerbl et al. 2023) requires slow, per-scene optimization. To address this, feed-forward models have emerged to perform 3D reconstruction from sparse or even single views in a single pass (Szymanowicz, Rupprecht, and Vedaldi 2024; Tang et al. 2024; Xu et al. 2024; Zhang et al. 2024b; Wu et al. 2024). Models like Splatter Image (Szymanowicz, Rupprecht, and Vedaldi 2024) and LGM (Tang et al. 2024) use U-Net-based architectures to directly regress Gaussian parameters from input images. Other works focus on scene-level generation by aggregating features from multiple images (Charatan et al. 2024; Chen et al. 2024; Hong et al. 2023; Li et al. 2023; Wang et al. 2023b).
@@
Figure 2: Architecture of PhysGM. The model conditions on one or more input views and their corresponding camera parameters, which are processed by a U-Net encoder to produce a shared latent representation $\mathbf{z}$. This latent is then decoded by two parallel heads: (1) a Gaussian Head predicting the initial 3D Gaussian scene parameters $\boldsymbol{\psi}$, and (2) a Physics Head that predicts a distribution over the object’s physical properties $\boldsymbol{\theta}$. The sampled parameters ($\boldsymbol{\psi},\boldsymbol{\theta}$) initialize a Material Point Method (MPM) simulator to generate the final dynamic sequence. The entire architecture is trained in a two-stage paradigm: first, supervised pre-training on ground-truth data establishes a robust generative prior. Subsequently, a DPO-based fine-tuning stage uses the ranks against a ground-truth video and aligns the model with physically plausible results.
Physics-Grounded 4D Synthesis
@@
The integration of physics into 4D generation is a nascent but critical field. The pioneering work, PhysGaussian (Xie et al. 2024), first coupled 3DGS with continuum mechanics simulation via the Material Point Method (MPM) in the Warp framework (Macklin 2022). This established a foundational pipeline but required expert-driven, manual configuration of physical parameters for each scene.
@@
Subsequent research has focused on automating this parameterization. One popular approach uses SDS to distill motion priors from video models, enabling data-driven optimization of material properties (Zhang et al. 2024c; Huang et al. 2025; Liu et al. 2024a; Tan et al. 2024). For instance, DreamPhysics (Huang et al. 2025) proposed a feature-disentanglement framework to separate motion and appearance. Others have explored using Large Language Models (LLMs) to infer parameters from text descriptions (Zhao et al. 2024; Chen et al. 2025; Liu et al. 2024b; Cai et al. 2024), though these can struggle with continuous physical quantities. More advanced methods learn neural constitutive models to represent complex material behaviors (Lin et al. 2025). Concurrently, efforts have been made to construct dedicated physical 3D datasets (Cao et al. 2025).
Despite these advances, the entire field suffers from two critical bottlenecks. First, these methods presuppose the existence of a high-quality, pre-optimized 3DGS model, requiring dense multi-view inputs and significant offline processing. Second, the process of imbuing this static model with physics is treated as a separate, per-scene optimization problem. This iterative fine-tuning, often taking hours to converge for a single object, renders existing approaches impractical for scalable content creation or any application requiring rapid generation. Our work, PhysGM, directly addresses these gaps. By integrating a feed-forward generative architecture with a learned physical prior, we eliminate the need for both per-scene optimization and pre-reconstructed 3D models. Our two-stage training paradigm allows us to produce physically grounded 4D simulations from sparse inputs at a speed several orders of magnitude faster than existing methods, bridging the gap between efficient generation and physical realism.
Method
Given a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$, our objective is to infer a physically-enabled 3D representation of the depicted object. This representation consists of two key components: (1) a set of 3D Gaussians, $\psi=\{g_{i}\}_{i=1}^{N}$ that defines the object’s geometry and appearance, and (2) a corresponding vector of physical properties $\theta$. Formally, we aim to learn a feed-forward mapping function $F$ such that: $F(I)=(\psi,\theta)$. This direct inference approach is designed to circumvent two critical bottlenecks in prior work: the need for slow, per-scene optimization to determine $\psi$, and the reliance on either manually defined physical priors $\theta_{\text{manual}}$ or unstable, heavy optimization to estimate $\theta$ from SDS loss. Our goal is to make both $\psi$ and $\theta$ predictable in a single feed-forward pass, enabling immediate, physics-based simulation. An overview of our pipeline is illustrated in Figure 2.
3D Gaussian Splatting
@@
The core of 3DGS (Kerbl et al. 2023) is to model the scene as a set of $N$ anisotropic 3D Gaussians. Each Gaussian is parameterized by: position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, defining its center in world space; covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ defining its shape and orientation; opacity $\alpha\in[0,1]$ controlling its transparency; color represented by Spherical Harmonics (SH) coefficients $\mathbf{c}$ to model view-dependent effects. The covariance matrix can be computed with a rotation matrix $\mathbf{R}$ and a diagonal scaling matrix $\mathbf{S}$: $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$
@@
During rendering, these 3D Gaussians are projected onto the 2D image plane for a given camera view. The final color $\mathbf{C}$ for a pixel is computed by alpha-blending the $N$ Gaussians that overlap the pixel:
$\mathbf{C}=\sum_{i=1}^{N}\mathbf{c}_{i}^{\prime}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}),$ (1)
where $\mathbf{c}_{i}^{\prime}$ is the view-dependent color evaluated from the SH coefficients, and $\alpha_{i}^{\prime}$ is the effective opacity of the $i$-th Gaussian. The objective of all parameters $\psi=\{(\boldsymbol{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\alpha_{i},\mathbf{c}_{i})\}_{i=1}^{N}$ is to minimize the difference between the rendered and ground-truth images.
Probabilistic Modeling
To infer the physical and appearance parameters of an object, we first derive a 3D-aware latent representation $\mathbf{z}$ from the single input image $I$. This is achieved in two steps. First, we use MVDream (Shi et al. 2023) to generate four consistent novel views from the input image. Second, a U-Net encoder fuses these views and their associated camera embeddings into the latent vector $\mathbf{z}$. This vector then conditions the subsequent prediction of physical properties and dynamic appearance parameters.
Probabilistic Physics Modeling

### v2

@@ Abstract
Despite advances in physics-based 3D motion synthesis, current methods face key limitations: reliance on pre-reconstructed 3D Gaussian Splatting (3DGS) built from dense multi-view images with time-consuming per-scene optimization; physics integration via either inflexible, hand-specified attributes or unstable, optimization-heavy guidance from video models using Score Distillation Sampling (SDS); and naïve concatenation of prebuilt 3DGS with physics modules, which ignores physical information embedded in appearance and yields suboptimal performance. To address these issues, we propose PhysGM, a feed-forward framework that jointly predicts 3D Gaussian representation and physical properties from a single image, enabling immediate simulation and high-fidelity 4D rendering. Unlike slow appearance-agnostic optimization methods, we first pre-train a physics-aware reconstruction model that directly infers both Gaussian and physical parameters. We further refine the model with Direct Preference Optimization (DPO), aligning simulations with the physically plausible reference videos and avoiding the high-cost SDS optimization. To address the absence of a supporting dataset for this task, we propose PhysAssets, a dataset of 50K+ 3D assets annotated with physical properties and corresponding reference videos. Experiments show that PhysGM produces high-fidelity 4D simulations from a single image in one minute, achieving a significant speedup over prior work while delivering realistic renderings.Our project page is at:https://hihixiaolv.github.io/PhysGM.github.io/
@@ 1 Introduction
Recent advances in 3D representation, particularly 3D Gaussian Splatting (3DGS) [kerbl20233d], have revolutionized novel-view synthesis for static scenes. The next frontier is to imbue these representations with dynamic, physically plausible behavior, unlocking applications in virtual reality, robotics [lu2024manigaussian], and autonomous systems [zhou2024drivinggaussian]. However, creating high-fidelity, physics-based 4D content remains a significant challenge, often demanding computationally expensive, per-scene optimization pipelines that are ill-suited for real-time or large-scale deployment.
@@
Current paradigms for physics-based 4D synthesis are hampered by a fundamental bottleneck: a reliance on slow, iterative optimization. The typical workflow involves first reconstructing a 3DGS model from dense multi-view images, then manually specifying its physical properties (e.g., stiffness, mass) via configuration files [xie2024physgaussian], and finally running a physics simulation. This process is not only computationally prohibitive but also lacks scalability and generalization. While recent works have explored learning physical properties via Score Distillation Sampling (SDS) from video models [lin2025omniphysgs, zhang2024physdreamer], which rely on gradient backpropagation through a differentiable physics simulator, incurring heavy per-scene optimization and poor efficiency; hence, they do not resolve the core efficiency bottleneck. Moreover, the common practice of naively concatenating a prebuilt 3DGS with a physics module overlooks physical cues encoded in appearance, causing suboptimal performance.
@@
To face these limitations, we ask a fundamental question: Can we bypass per-scene optimization entirely and instead learn a generative model that produces a complete, physically-grounded 4D simulation in a single, feed-forward pass? This requires reframing the problem from one of slow, iterative reconstruction to one of amortized, feed-forward inference. To this end, we propose PhysGM, a feed-forward framework that enables optimization-free synthesis of dynamic 4D scenes from image inputs. Our key insight is a novel two-stage training paradigm designed to learn a generalizable physical prior and perceptual realism. In the first stage, we pre-train our model on a substantial dataset to jointly predict a 3D Gaussian representation and its corresponding physical properties—this joint optimization not only enhances the accuracy of both geometry and physics predictions but also establishes a well generative prior [wang2025vggt, dosovitskiy2020image]. Critically, it eliminates the reliance on pre-reconstructed 3DGS that demands multi-view images and time-consuming per-scene optimization, while addressing the flaw of overlooking visually embedded physical information caused by the naive concatenation of separate Gaussian and physics prediction modules. In the second stage, we employ Direct Preference Optimization (DPO) [rafailov2023direct] to fine-tune the model. By ranking generated simulations against ground-truth videos, we create preference pairs that effectively guide the model towards producing physically plausible and temporally coherent dynamic sequences-unlike SDS-based methods, this preference-driven approach completely eliminates the need for both time-consuming per-scene optimization processes and a differentiable physics engine.
@@
Given the lack of datasets pairing 3D assets with physical annotations and reference simulations for this task, we construct and release the PhysAssets Dataset, a substantial benchmark of over 50,000 3D assets, each annotated with its physical material properties and a corresponding physically-plausible simulation video, providing a critical resource for training and evaluating generative 4D models.
@@
  • We propose PhysGM, the first feed-forward framework capable of generating a physically-grounded 4D Gaussian simulation from an image input in one minute.
@@
  • We construct a novel two-stage training paradigm that combines large-scale supervised pre-training with DPO-based refinement, enabling the model to learn a physical prior and then align it with perceptual quality.
@@
  • We present the PhysAssets Dataset, a new substantial benchmark with 50,000+ 3D assets annotated with physical properties and simulation videos, to spur future research in this challenging domain.
  • Extensive experiments show that PhysGM produces high-fidelity 4D simulations from a single image in one minute, achieving substantial speedups over SDS-based baselines while delivering superior visual quality across metrics.
2 Related Work
2.1 4D Content Generation
Recent advances in 4D content generation—dynamic 3D scenes—primarily follow two paradigms. The first distills knowledge from pre-trained 2D/video diffusion models into a dynamic 3D representation using Score Distillation Sampling (SDS)[poole2022dreamfusion, yin20234dgen, ren2023dreamgaussian4d, ling2024align] or trajectory guidance [wang2025physctrl]. To ensure temporal coherence, methods employ explicit deformation models[wang2023modelscope] or multi-view video supervision [zhang20244diffusion, jiang2024animate3d]. However, the iterative nature of SDS is computationally intensive and prone to instability. Mitigation strategies include pre-synthesizing videos for direct supervision [ren2023dreamgaussian4d, pan2024efficient4d] or training generalizable models to bypass per-scene optimization [ren2024l4gm].
The second paradigm trains generative models directly on 3D data [jun2023shap, nichol2022point]. These models have evolved to generate diverse representations, from point clouds [nichol2022point] to implicit functions (SDFs)[jun2023shap] and explicit 3D Gaussians[tang2024lgm]. Its primary limitation is the scarcity of large-scale 3D datasets.
@@ 2 Related Work
Despite impressive visual results, both paradigms are fundamentally physics-agnostic. They produce temporally coherent but often physically implausible animations, as they lack an explicit model of real-world dynamics.
2.2 Feed-forward 3D Gaussian Models
@@
Parallel to 4D generation, another line of research has focused on accelerating the creation of static 3D scenes. Traditional 3DGS reconstruction [kerbl20233d] requires slow, per-scene optimization. To address this, feed-forward models have emerged to perform 3D reconstruction from sparse or even single views in a single pass [szymanowicz2024splatter, tang2024lgm, xu2024grm, zhang2024gs, wu2024direct3d, li2025lirm]. Models like Splatter Image [szymanowicz2024splatter] and LGM [tang2024lgm] use U-Net-based architectures to directly regress Gaussian parameters from input images, while LVSM [jin2024lvsm] outputs only rendered images instead of 3D Gaussians. Other works focus on scene-level generation by aggregating features from multiple images [charatan2024pixelsplat, chen2024mvsplat, hong2023lrm, li2023instant3d, wang2023pf, ziwen2025long].
@@
Figure 2: Pipeline of PhysGM. The model conditions on one or four input views and their corresponding camera parameters, which are processed by a transformer-based model to produce output tokens. These tokens then decoded by two parallel heads: (1) a DPT Head predicting the initial 3D Gaussian scene parameters $\boldsymbol{\psi}$, and (2) a Physics Head that predicts a distribution over the object’s physical properties $\boldsymbol{\theta}$. The sampled parameters ($\boldsymbol{\psi},\boldsymbol{\theta}$) initialize a Material Point Method (MPM) simulator to generate the final dynamic sequence. The entire architecture is trained in a two-stage paradigm: first, supervised pre-training on ground-truth data establishes a well generative prior. Subsequently, a DPO-based fine-tuning stage uses the ranks against a ground-truth video and aligns the model with physically plausible results.
2.3 Physics-Grounded 4D Synthesis
@@
The integration of physics into 4D generation is a nascent but critical field. The pioneering PhysGaussian [xie2024physgaussian] first coupled 3DGS with Material Point Method (MPM) simulations based on warp [macklin2022warp], but its reliance on manual, per-scene parameter tuning highlighted the need for automation.
@@
Subsequent research has focused on automating this process. A dominant approach uses Score Distillation Sampling (SDS) to distill motion priors from video models, enabling data-driven optimization of material properties [zhang2024physdreamer, huang2025dreamphysics, liu2024physics3d, tan2024physmotion]. Others leverage Large Language Models (LLMs) to infer parameters from text [zhao2024efficient, chen2025physgen3d, liu2024physgen, cai2024gic], though accurately predicting continuous values remains a challenge. More advanced methods learn neural constitutive models [lin2025omniphysgs].Concurrently, efforts have been made to construct dedicated physical 3D datasets [cao2025physx].
However, existing approaches are fundamentally limited by their reliance on pre-reconstructed models and slow, per-scene optimization. We propose PhysGM to overcome these bottlenecks. Our framework enables rapid, end-to-end 4D synthesis from sparse inputs by learning a physical prior, unifying efficiency with physical realism.
3 Method
We present PhysGM, a transformer-based physical reconstruction model that, given the posed RGB images $I=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{n}$, the model predicts in one forward pass: (1) a set of 3D Gaussians for geometry and appearance, and (2) a vector of physical properties for physical simulation. We first review the fundamentals of 3D Gaussian Splatting (Sec. 3.1). We then detail our transformer-based architecture that jointly predicts 3DGS and physics parameters (Sec. 3.2). Next, we explain how these parameters drive a Material Point Method (MPM) simulation (Sec. 3.3) and introduce our DPO-based strategy for refining the model with ground-truth dynamics (Sec. 3.4). Finally, we present PhysAssets, a new dataset created to train and evaluate our model (Sec. 3.5).
3.1 3D Gaussian Splatting
@@ 3 Method
The core of 3DGS [kerbl20233d] is to model the scene as a set of $N$ anisotropic 3D Gaussians. Each Gaussian is parameterized by: position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, defining its center in world space; covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ defining its shape and orientation; opacity $\alpha\in[0,1]$ controlling its transparency; color represented by Spherical Harmonics (SH) coefficients $\mathbf{c}$ that enable modeling of view-dependent appearance effects.
@@
The covariance matrix can be computed with a rotation matrix $\mathbf{R_{\text{mat}}}$ and a diagonal scaling matrix $\mathbf{S}$: $\boldsymbol{\Sigma}=\mathbf{R_{\text{mat}}}\mathbf{S}\mathbf{S}^{T}\mathbf{R_{\text{mat}}}^{T}$, where $\mathbf{S}=diag(\mathbf{s})$. During rendering, these 3D Gaussians are projected onto the 2D image plane for a given camera view. The final color $\mathbf{C}$ for a pixel is computed by alpha-blending the $N$ Gaussians that overlap the pixel with its $\alpha_{i}$ and color $\mathbf{c}_{i}$ . The objective of all parameters $\psi=\{(\boldsymbol{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\alpha_{i},\mathbf{c}_{i})\}_{i=1}^{N}$ is to minimize the difference between the rendered and ground-truth images.
3.2 Model Architecture
We propose a transformer-based architecture PhysGM, designed to jointly regress 3DGS parameters and physical attributes from posed images. Our framework comprises three main components: multi-modality encoding, a transformer-based backbone, and decoders for 3DGS and physics properties. An overview is shown in Figure 2.
Multi-modality tokenization
We adopt DINOv3 [siméoni2025dinov3] as image encoder $E_{\text{img}}$. Each input image $I_{i}$ is patchified into tokens, which are then projected by a linear layer to produce high-dimensional image features. To explicitly encode camera geometry, we represent the principal ray of each pixel for view $I_{i}$ using Plücker ray coordinates $C_{i}$. We process $C_{i}$ by a dense representation encoder [zhang2025ufmsimplepathunified] $E_{\text{cam}}$ to produce geometry tokens. The resulting camera tokens are then concatenated with corresponding image tokens. The concatenated tokens ${(\textbf{t}_{i})}_{i=1}^{N}$ are formed by concatenating the corresponding image and camera tokens. We further prepend three learnable global tokens $\mathbf{g}_{1},\mathbf{g}_{2},\mathbf{g}_{3}$} to the sequence. These tokens aggregate global scene information and will be used for physics prediction. The concatenated tokens $\textbf{t}_{i}$ and input tokens $\mathcal{T}_{\text{in}}$ are thus formed as:
$\textbf{t}_{i}=\text{concat}(E_{\text{img}}(I_{i}),E_{\text{cam}}(C_{i})),$ (1)
$\mathcal{T}_{\text{in}}={(\textbf{t}_{i})}_{i=1}^{N}\cup\mathbf{g}_{k},k=1,2,3.$ (2)
During inference with a single input image, we use MVAdapter [huang2025mv] to synthesize three fixed auxiliary views—rear, left, and right—while the original image serves as the frontal view.
Transformer Backbone
The complete sequence of input tokens $\mathcal{T}_{\text{in}}$ is fed into our 24-layer transformer backbone to learn the contextualized representations. We collect the output tokens from the intermediate layers, which provide a multi-scale representation of the scene, capturing both high-level semantics and low-level details.
Prediction heads
We employ two distinct, specialized heads to map the backbone’s output tokens to 3DGS parameters and physical attributes.
To predict the 3DGS representation, we adopt a Dense Prediction Transformer (DPT) [ranftl2021vision, yang2024depth] head $f_{\text{gs}}$. This head takes the multi-scale features from the backbone and progressively upsamples them through a series of refinement stages. For each input view, it outputs per-pixel maps corresponding to the 3DGS parameters $\psi_{i}=f_{\text{gs}}(\textbf{t}_{i})$. The Gaussians predicted from all views are then aggregated to form the final, coherent 3D scene representation. During pre-training, we optimize the network by jointly minimizing the MSE, alpha and LPIPS [zhang2018unreasonable] losses between rendered images and the ground-truth views.